A rectangle is drawn so that the width is 4 feet shorter than the length. The area of the rectangle is 12 square feet. Find the length of the rectangle.

Answer by ewatrrr(24785)

 
Hi,
.
L(L-4) = 12ft^2
L^2 - 4L - 12 = 0 (Tossing out the negative solution for unit measure)
(L+2)(L-6) = 0
L = 6ft
And...checking
(6ft)(2ft) = 12ft^2